OpenStudy (anonymous):

Complex Numbers!!! >_< Simplify 7/10i

OpenStudy (amistre64):

i dunno, that looks pretty simple already

OpenStudy (amistre64):

usually this tends to mean multiply by the conjugate ....

OpenStudy (zarkon):

I assume you mean \(\displaystyle\frac{7}{10i}\)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

All you need - is multiply by i and divide by i, and remember, that i*i = -1, so... 7/(10i) = (7i)/(10i*i) = -(7i)/10 = -(7/10)*i

OpenStudy (anonymous):

K. Thank you!